# OmniJARVIS: Unified Vision-Language-Action Tokenization Enables Open-World Instruction Following Agents

### v1

@@ Abstract
This paper presents OmniJARVIS, a novel Vision-Language-Action (VLA) model for open-world instruction-following agents in Minecraft. Compared to prior works that either emit textual goals to separate controllers or produce the control command directly, OmniJARVIS seeks a different path to ensure both strong reasoning and efficient decision-making capabilities via unified tokenization of multimodal interaction data. First, we introduce a self-supervised approach to learn a behavior encoder that produces discretized tokens for behavior trajectories $\tau=\{o_{0},a_{0},\dots\}$ and an imitation learning (IL) policy decoder conditioned on these tokens. These additional behavior tokens will be augmented to the vocabulary of pretrained Multimodal Language Models (MLMs). With this encoder, we then pack long-term multimodal interactions involving task instructions, memories, thoughts, observations, textual responses, behavior trajectories, etc. into unified token sequences and model them with autoregressive transformers. Thanks to the semantically meaningful behavior tokens, the resulting VLA model, OmniJARVIS, can reason (by producing chain-of-thoughts), plan, answer questions, and act (by producing behavior tokens for the IL policy decoder). OmniJARVIS demonstrates excellent performances on a comprehensive collection of atomic, programmatic, and open-ended tasks in open-world Minecraft. Our analysis further unveils the crucial design principles in interaction data formation, unified tokenization, and its scaling potentials. The dataset, models, and code will be released at https://craftjarvis.org/OmniJarvis/.
[…]
Yitao Liang
Zihao Wang<zhwang@stu.pku.edu.cn>, Shaofei Cai<caishaofei@stu.pku.edu.cn>, Anji Liu<liuanji@cs.ucla.edu>, Xiaojian Ma<xiaojian.ma@ucla.edu>, Yitao Liang<yitaol@pku.edu.cn>
@@ 1 Introduction
Figure 1: Illustration of multi-modal interaction data for decision-making. A canonical interaction sequence depicting the human decision-making process starts from a given task instruction and memory, followed by a series of sub-task completion which involves initial observations, chain-of-thought reasoning, and behavior trajectories. Our proposed VLA model OmniJARVIS jointly models the vision (observations), language (instructions, memories, thoughts), and actions (behavior trajectories) as unified autoregressive sequence prediction. A self-supervised behavior encoder (detailed in Section 2 and Figure 2) converts the actions into behavior tokens while the other modalities are tokenized following the practices of MLMs (Liu et al., 2024; Bavishi et al., 2023; Alayrac et al., 2022).
@@
Upon the success of pretrained Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023; Gong et al., 2023; Cheng et al., 2024) and Multimodal Langauge Models (MLMs) (Liu et al., 2024; Hinck et al., 2024; Alayrac et al., 2022; Zhao et al., 2023; Man et al., 2024), some recent works have been venturing into developing Vision-Language-Action (VLA) models (Brohan et al., 2023; Huang et al., 2023; Wang et al., 2023b; Reed et al., 2022), a promising pathway towards the ultimate goal of building autonomous agents that can follow and even self-generated instructions to fulfill various reasoning and acting tasks in open world environments. Among them, two most prominent architectures have been proposed: 1) Combining an off-the-shelf MLM (Liu et al., 2024; Alayrac et al., 2022) with separate goal-conditioned controllers (Lifshitz et al., 2023; Cai et al., 2023b, a), where MLM reasons, plans and pilots the controllers by producing textual goal instructions, e.g. DEPS (Wang et al., 2023c), JARVIS-1 (Wang et al., 2023b), voyager (Wang et al., 2023a); 2) Tuning a pretrained MLM into producing control commands directly, while maintaining the reasoning and language capabilities, e.g. RT-2 (Brohan et al., 2023), LEO (Huang et al., 2023). However, these two designs could still have significant drawbacks when it comes to open-world environments. First, an open world (e.g., Minecraft) usually teams up with an infinite number of complex and highly contextualized tasks (Fan et al., 2022; Lin et al., 2023), and it can be fairly challenging to depict them in text only. Therefore, VLA models that solely depend on text to communicate with the text-conditioned policies (Wang et al., 2023b, c) may fail to correctly pilot these controllers. On the other side, emitting the control command directly (Brohan et al., 2023; Huang et al., 2023) without invoking separate controllers could alleviate the aforementioned communication problem but given the long-horizon nature of open-world tasks, it is less practical to perform long-term control with a large VLA model as the context length requirement, computation cost and inference efficiency could become unaffordable.
@@
In this paper, we aim to tackle the aforementioned issues of existing VLA models when facing open-world environments: complex & context-dependent tasks and long-term tasks. Our key insight originates from the observation of human decision-making: Given these open-world tasks, humans can make informed decisions via multi-round mental, verbal, and physical interactions (an illustration can be found in Figure 1). Therefore, if the VLA model can manage to learn from such interaction data, it may master the underlying human decision-making procedures. However, modeling interaction data is non-trivial: it is multi-modal, encloses vision (mostly observations), language (instructions, thoughts, etc.), and actions (behavior trajectories). Compared to the fruitful explorations on jointly tokenizing vision and language (Liu et al., 2024; Bavishi et al., 2023; Vinyals et al., 2015; Alayrac et al., 2022) into sequences for autoregressive modeling (Brown et al., 2020), tokenizing behavior trajectories (actions) is hard due to the following reasons. On the one hand, directly using low-level actions from the environment would pose huge challenges to the model’s ability to process long sequences, which significantly hurts performance. It also hinders us from leveraging the planning ability of generative models. On the other hand, language-level action tokens require significantly more supervision and cannot accurately describe all possible actions.
Figure 2: Self-supervised learning for behavior tokenizer of OmniJARVIS. We modify the VAE-based self-supervised learning of behavior trajectories in (Cai et al., 2023b) to train the behavior tokenizer and de-tokenizer in OmniJARVIS. Specifically, we adopt the auto-encoding objective but replace the Gaussian latent with a discrete representation based on Finite Scalar Quantizer (Mentzer et al., 2023). The encoder will then be used as the behavior tokenizer to produce discrete tokens from the actions (behavior trajectories) in multimodal interaction data, while the behavior tokens emitted by OmniJARVIS will be sent to the policy decoder to perform motor control.
@@
To this end, we propose OmniJARVIS, a novel VLA model that jointly models vision, language, and actions in interaction data with unified tokenization. OmniJARVIS comprises two key ideas: 1) Behavior Tokenization. We introduce a self-supervised approach to learn a behavior encoder that produces discretized tokens for actions (behavior trajectories) and an imitation learning policy decoder conditioned on these tokens (Section 2); 2) Autoregressive Modeling. By augmenting these behavior tokens into the vocabulary of pretrained MLMs, we pack the multimodal interaction data into unified token sequences and learn a transformer on these sequences with an autoregressive modeling objective. We conduct comprehensive evaluations in the open-world Minecraft Universe (Lin et al., 2023). OmniJARVIS demonstrates impressive performances on a wide range of atomic, programmatic, and open-ended Minecraft tasks. Our analysis confirms several critical design choices in data formation, tokenization, and the scaling potential of OmniJARVIS. Our contributions are as follows:
@@
  • We conduct extensive evaluations in open-world Minecraft Universe to demonstrate OmniJARVIS’s proficiency across various tasks, and present in-depth analyses to reveal valuable insights.
@@ 2 A Tokenizer for Behaviors
Specifically, we aim at producing a set of $N$ discrete behavior tokens $s^{\text{bhv}}_{1},\dots,s^{\text{bhv}}_{N}$ from a behavior trajectory $\tau=\{o_{0},a_{0},\dots\}$. Further, a de-tokenizer is needed to map these tokens back to an action rollout in the environment that reproduces the goal achieved in $\tau$. GROOT (Cai et al., 2023b) explores a VAE-based approach to jointly learn a latent representation of behavior trajectories and an imitation learning policy decoder that conditions the latent as goal. However, the continuous latent cannot be used as the behavior tokens as they can be more difficult to learn and decode with the existing discrete tokens of pretrained MLMs (Huang et al., 2023; Ma et al., 2022). Therefore, we replace the Gaussian latent in GROOT with an improved vector quantized discrete latent called Finite Scalar Quantization (FSQ) (Mentzer et al., 2023). We adopt a quantization configuration of $[8,8,8,6,5]$, which means a code with a length=5 and a codebook size of $8\times 8\times 8\times 6\times 5=15360$ is produced. The configuration is selected by a simple grid search. Overall, the behavior tokenizer (behavior encoder) $e_{\phi}(o_{1;T})$ and the de-tokenizer (IL policy decoder) $\pi_{\theta}(a_{t}|o_{1:t})$ is learned with the following objective:
$\underset{(\phi,\theta)}{\text{argmin}}\leavevmode\nobreak\ \mathbb{E}_{\tau\sim\mathcal{D}}\left[\sum_{t=1}^{T}-\log\pi_{\theta}(a_{t}|o_{1:t},f(e_{\phi}(o_{1:T})))\right],$ (1)
@@
Compared to our behavior tokenization, most prior work in VLA models, either represents the behavior trajectories in interaction data as a textual goal description and invokes a separate goal-conditioned controller (Wang et al., 2023b, c), or represents the state-action sequence $\{o_{0},a_{0},\dots\}$ directly as in Decision Transformers (DT) (Chen et al., 2021; Huang et al., 2023; Reed et al., 2022; Brohan et al., 2023). Our approach offers a more compact but still informative representation of the actions part in multimodal interaction data. Moreover, the action readout, i.e. simply sending the behavior tokens to the policy decoder, is also more efficient than the DT-style direct control from VLA models (Reed et al., 2022; Brohan et al., 2023; Huang et al., 2023).
@@ 3 Multimodal Interaction Data and OmniJARVIS
As illustrated in Figure 1, canonical multimodal interaction data comprises vision (observations), language (instructions, memories, thoughts), and actions (behavior trajectories). However, it can be difficult to directly collect such interaction data from human annotators. Therefore, we propose to convert an existing Minecraft gameplay dataset (Baker et al., 2022) into the multimodal interaction data required by OmniJARVIS. We begin with a formal definition of the interaction data, followed by our approach for data conversion and augmentation from existing datasets, and finish up with the architecture, formulation of learning on such interaction data, and inference procedure of OmniJARVIS. An overview of OmniJARVIS architecture and inference can be found in Figure 3.
@@ 3.2 Preparing Multimodal Interaction Data
In reality, many segments of the multimodal interaction $\mathbb{D}$ can be missing in public datasets. We consider the Minecraft contractor data released by OpenAI (Baker et al., 2022) and it only contains behavior trajectories $D^{\text{bhv}}_{t}$. Therefore, we need to properly augment the data with the additional textual segments including instructions $D^{\text{inst}}_{t}$, memory $D^{\text{mem}}_{t}$, and thoughts $D^{\text{tht}}_{t}$. We follow the prior practices (Huang et al., 2023; Liu et al., 2024) to synthesize the required text using LLMs. Below, we detail how each type of segment is constructed. More details can be found in appendix.
@@
The instruction is a high-level description of what task is being performed in the current interaction sequence. The considered OpenAI Minecraft data includes meta information of each gameplay video, which depicts fundamental events that happened during in Minecraft gameplay, e.g. what block was just destroyed, what entity was just killed, what item was just crafted, etc. Such meta-information can provide a basic overview of what the player has been through in the gameplay. We therefore prompt an LLM into summarizing the gameplay with the meta information. The summary will then be used as the instruction $D^{\text{inst}}_{t}$ of the current interaction trajectory.
@@
The memory is the summary of what agents have finished in the previous interaction sequences. Due to the limited sequence length that the auto-regressive model can handle, the model needs to learn to summarize key information related to the task in historical interactions and ignore behaviors unrelated to instructions. The memory will be updated based on the results of each episode trunk and used for subsequent episode trunks. We therefore prompt an LLM into summarizing the gameplay with the meta information. The summary will then be used as the memory $D^{\text{mem}}_{t}$ of the current interaction trajectory. The memory prompt can be found in Appendix G.
@@
The thought is the agent’s reasoning and explanation of its own decisions. Previous methods have confirmed that using thought-enhanced interaction data helps language models understand decision-making (Hu and Clune, 2024). Compared to labeling thoughts by humans (Yao et al., 2022; Brohan et al., 2023), we assume that thought is an intermediate variable that can be determined by the actions taken and observations made before and after the action, which is similar to an Inverse Dynamics Model (Baker et al., 2022). We therefore prompt an LLM into estimating the thought of decisions with in-context learning, which will then be used as the thought $D^{\text{tht}}_{t}$ of the current behavior. Details can be found in Appendix F.
@@ 3.3 Architecture, Training, and Inference of OmniJARVIS
As illustrated in Figure 3, OmniJARVIS is built upon a pretrained MLM. We augment the original vocabulary of the MLM with additional tokens from the behavior tokenizer. Specifically, as we adopted the $[8,8,8,6,5]$ FSQ configuration (Section 2), we augment with $8+8+8+6+5=35$ new tokens as each behavior comprises 5 behavior tokens $s^{bhv}_{1},\dots,s^{bhv}_{5}$ corresponding to 5 FSQ level. We formulate the learning objective of OmniJARVIS following (Brown et al., 2020; Raffel et al., 2020) in a prefix language modeling fashion. For a batch $\mathcal{B}$ of token sequence $s$, we optimize OmniJARVIS via:
@@ 4.1 Overview
The training of the OmniJARVIS is divided into two stages. In the first step, we use a self-supervised training method to train a Behavior Tokenizer, including the Encoder and Decoder jointly. We use FSQ as a quantization method and build a codebook with 8*8*8*6*5 discrete codes. The training data for Behavior Tokenizer comes from Contractor Dataset (Baker et al., 2022), which is a collection of Minecraft gameplay videos. The training parameters and details remain consistent with GROOT, which can be found in Appendix A.
In the second stage, we use this behavior tokenizer to process Minecraft offline trajectories to obtain behavior token sequences. We add 35 (8+8+8+6+5) additional tokens to the MLM tokenizer as behavior tokens for unified representation, so each time the VLA needs to output a continuous sequence of 5 tokens to represent a complete behavior. We use GPT-3.5 to synthesize thought, memory, and instruction to raw offline datasets to build complete interaction data. The specific prompt can be found in Appendix F. These data collectively constitute the embodied instruction-following dataset of OmniJARVIS, including 600k trajectories and about 900M tokens.
The training dataset of OmniJARVIS further includes a large amount of QA data about Minecraft. We generate a large number of seed questions about these texts using web pages on the Minecraft wiki. Then, we use the self-instruct method to generate a large number of creative questions and instructions. This constructed QA dataset consists of 300k conversations with about 90M tokens. During the training process, the QA data and instruction-following data are mixed, with a total of about 1T tokens, to train OmniJARVIS. In specific, we SFT (supervised finetune) LLaVA-7B (Liu et al., 2024). The details can be found in Appendix A. To further demonstrate the generalizability of the method, we also fine-tune LLaVA at different scales and VLM Fuyu-8B with different architectures. The relevant results are presented in Section 4.5 and Section 4.6.
@@
We conduct experiments in the complex and open-world environment of Minecraft, a voxel-based 3D video game that has garnered significant attention from real-life research due to its popularity and diverse mechanics (Guss et al., 2019; Fan et al., 2022). We first evaluate OmniJARVIS with atomic tasks, which are skill-level tasks, testing VLAs’ ability to follow simple and straightforward instructions. Then we evaluate OmniJARVIS with programmatic tasks, which require the agent to obtain an item starting from an empty inventory. The success of these tasks requires VLAs to decompose the provided instruction into atomic-level subtasks, and hence tests VLAs’ complex reasoning ability. Finally, we test OmniJARVIS with open-world embodied question-answering benchmarks and creative free-form instruction-following. We also conduct ablation experiments of OmniJARVIS with different behavior tokenizers, different training dataset formats, and different vision tokenizations. Finally, we explore the scaling potential of OmniJARVIS with different models and data scales.
4.2 Main Results I: Atomic Tasks

### v2

# OmniJARVIS Unified Vision-Language-Action Tokenization Enables Open-World Instruction Following Agents

@@ Abstract
This paper presents OmniJARVIS, a novel Vision-Language-Action (VLA) model for open-world instruction-following agents in Minecraft. Compared to prior works that either emit textual goals to separate controllers or produce the control command directly, OmniJARVIS seeks a different path to ensure both strong reasoning and efficient decision-making capabilities via unified tokenization of multimodal interaction data. First, we introduce a self-supervised approach to learn a behavior encoder that produces discretized tokens for behavior trajectories $\tau=\{o_{0},a_{0},\dots\}$ and an imitation learning policy decoder conditioned on these tokens. These additional behavior tokens will be augmented to the vocabulary of pretrained Multimodal Language Models. With this encoder, we then pack long-term multimodal interactions involving task instructions, memories, thoughts, observations, textual responses, behavior trajectories, etc. into unified token sequences and model them with autoregressive transformers. Thanks to the semantically meaningful behavior tokens, the resulting VLA model, OmniJARVIS, can reason (by producing chain-of-thoughts), plan, answer questions, and act (by producing behavior tokens for the imitation learning policy decoder). OmniJARVIS demonstrates excellent performances on a comprehensive collection of atomic, programmatic, and open-ended tasks in open-world Minecraft. Our analysis further unveils the crucial design principles in interaction data formation, unified tokenization, and its scaling potentials. The dataset, models, and code will be released at https://craftjarvis.org/OmniJARVIS/.
@@ 1 Introduction
Upon the success of pretrained Large Language Models (LLMs) [7, 35, 40, 17, 13] and Multimodal Langauge Models (MLMs) [31, 20, 1, 53, 33], some recent works have been venturing into developing Vision-Language-Action (VLA) models [6, 22, 47, 38], a promising pathway towards the ultimate goal of building autonomous agents that can follow and even self-generated instructions to fulfill various reasoning and acting tasks in open world environments. Among them, two most prominent architectures have been proposed: 1) Combining an off-the-shelf MLM [31, 1] with separate goal-conditioned controllers [28, 10, 9], where MLM reasons, plans and pilots the controllers by producing textual goal instructions, e.g. DEPS [46], JARVIS-1 [47], voyager [44]; 2) Tuning a pretrained MLM into producing control commands directly, while maintaining the reasoning and language capabilities, e.g. RT-2 [6], LEO [22]. However, these two designs could still have significant drawbacks when it comes to open-world environments. First, an open world (e.g., Minecraft) usually teams up with an infinite number of complex and highly contextualized tasks [16, 29], and it can be fairly challenging to depict them in text only. Therefore, VLA models that solely depend on text to communicate with the text-conditioned policies [47, 46] may fail to correctly pilot these controllers. On the other side, emitting the control command directly [6, 22] without invoking separate controllers could alleviate the aforementioned communication problem but given the long-horizon nature of open-world tasks, it is less practical to perform long-term control with a large VLA model as the context length requirement, computation cost and inference efficiency could become unaffordable.
@@
In this paper, we aim to tackle the aforementioned issues of existing VLA models when facing open-world environments: complex & context-dependent tasks and long-term tasks. Our key insight originates from the observation of human decision-making: Given these open-world tasks, humans can make informed decisions via multi-round mental, verbal, and physical interactions (an illustration can be found in Figure 1). Therefore, if the VLA model can manage to learn from such interaction data, it may master the underlying human decision-making procedures. However, modeling interaction data is non-trivial: it is multi-modal, encloses vision (mostly observations), language (instructions, thoughts, etc.), and actions (behavior trajectories). Compared to the fruitful explorations on jointly tokenizing vision and language [31, 3, 43, 1] into sequences for autoregressive modeling [7], tokenizing behavior trajectories (actions) is hard due to the following reasons. On the one hand, directly using low-level actions from the environment would pose huge challenges to the model’s ability to process long sequences, which significantly hurts performance. It also hinders us from leveraging the planning ability of generative models. On the other hand, language-level action tokens require significantly more supervision and cannot accurately describe all possible actions.
Figure 1: Illustration of multi-modal interaction data for decision-making. A canonical interaction sequence depicting the human decision-making process starts from a given task instruction and memory, followed by a series of sub-task completion which involves initial observations, chain-of-thought reasoning, and behavior trajectories. Our proposed VLA model OmniJARVIS jointly models the vision (observations), language (instructions, memories, thoughts), and actions (behavior trajectories) as unified autoregressive sequence prediction. A self-supervised behavior encoder (detailed in Section 2 and Figure 2) converts the actions into behavior tokens while the other modalities are tokenized following the practices of MLMs [31, 3, 1].
@@
To this end, we propose OmniJARVIS, a novel VLA model that jointly models vision, language, and actions in interaction data with unified tokenization. OmniJARVIS comprises two key ideas: 1) Behavior Tokenization. We introduce a self-supervised approach to learn a behavior encoder that produces discretized tokens for actions (behavior trajectories) and an imitation learning policy decoder conditioned on these tokens (Section 2); 2) Autoregressive Modeling. By augmenting these behavior tokens into the vocabulary of pretrained MLMs, we pack the multimodal interaction data into unified token sequences and learn a transformer on these sequences with an autoregressive modeling objective. We conduct comprehensive evaluations in the open-world Minecraft Universe [29]. OmniJARVIS demonstrates impressive performances on a wide range of atomic, programmatic, and open-ended Minecraft tasks. Our analysis confirms several critical design choices in data formation, tokenization, and the scaling potential of OmniJARVIS. Our contributions are as follows:
@@
  • We conduct extensive evaluations in open-world Minecraft to demonstrate OmniJARVIS’s proficiency across various tasks and present in-depth analyses to reveal valuable insights.
Figure 2: Self-supervised learning for behavior tokenizer of OmniJARVIS. We modify the VAE-based self-supervised learning of behavior trajectories in [10] to train the behavior tokenizer and de-tokenizer in OmniJARVIS. Specifically, we adopt the auto-encoding objective but replace the Gaussian latent with a discrete representation based on Finite Scalar Quantizer [34]. The encoder will then be used as the behavior tokenizer to produce discrete tokens from the actions (behavior trajectories) in multimodal interaction data, while the behavior tokens emitted by OmniJARVIS will be sent to the policy decoder to perform motor control.
@@ 2 A Tokenizer for Behaviors
Specifically, we aim at producing a set of $N$ discrete behavior tokens $s^{\text{bhv}}_{1},\dots,s^{\text{bhv}}_{N}$ from a behavior trajectory $\tau=\{o_{0},a_{0},\dots\}$. Further, a de-tokenizer is needed to map these tokens back to an action rollout in the environment that reproduces the goal achieved in $\tau$. GROOT [10] explores a VAE-based approach to jointly learn a latent representation of behavior trajectories and an imitation learning policy decoder that conditions the latent as goal. However, the continuous latent cannot be used as the behavior tokens as they can be more difficult to learn and decode with the existing discrete tokens of pretrained MLMs [22, 32]. Therefore, we replace the Gaussian latent in GROOT with an improved vector quantized discrete latent called Finite Scalar Quantization (FSQ) [34]. We adopt a quantization configuration of $[8,8,8,6,5]$, which means a code with a length=5 and a codebook size of $8\times 8\times 8\times 6\times 5=15360$ is produced. The configuration is selected by a simple grid search. Overall, the behavior tokenizer (behavior encoder) $e_{\phi}(o_{1;T})$ and the de-tokenizer (IL policy decoder) $\pi_{\theta}(a_{t}|o_{1:t})$ is learned with the following objective:
$\underset{(\phi,\theta)}{\text{argmin}}~{}\mathbb{E}_{\tau\sim\mathcal{D}}\left[\sum_{t=1}^{T}-\log\pi_{\theta}(a_{t}|o_{1:t},f(e_{\phi}(o_{1:T})))\right],$ (1)
@@
Compared to our behavior tokenization, most prior work in VLA models, either represents the behavior trajectories in interaction data as a textual goal description and invokes a separate goal-conditioned controller [47, 46], or represents the state-action sequence $\{o_{0},a_{0},\dots\}$ directly as in Decision Transformers (DT) [11, 22, 38, 6]. Our approach offers a more compact but still informative representation of the actions part in multimodal interaction data. Moreover, the action readout, i.e. simply sending the behavior tokens to the policy decoder, is also more efficient than the DT-style direct control from VLA models [38, 6, 22].
@@ 3 Multimodal Interaction Data and OmniJARVIS
As illustrated in Figure 1, canonical multimodal interaction data comprises vision (observations), language (instructions, memories, thoughts), and actions (behavior trajectories). However, it can be difficult to directly collect such interaction data from human annotators. Therefore, we propose to convert an existing Minecraft gameplay dataset [2] into the multimodal interaction data required by OmniJARVIS. We begin with a formal definition of the interaction data, followed by our approach for data conversion and augmentation from existing datasets, and finish up with the architecture, formulation of learning on such interaction data, and inference procedure of OmniJARVIS. An overview of OmniJARVIS architecture and inference can be found in Figure 3.
@@ 3.2 Preparing Multimodal Interaction Data
In reality, many segments of the multimodal interaction $\mathbb{D}$ can be missing in public datasets. We consider the Minecraft contractor data released by OpenAI [2] and it only contains behavior trajectories $D^{\text{bhv}}_{t}$. Therefore, we need to properly augment the data with the additional textual segments including instructions $D^{\text{inst}}_{t}$, memory $D^{\text{mem}}_{t}$, and thoughts $D^{\text{tht}}_{t}$. We follow the prior practices [22, 31] to synthesize the required text using LLMs. Below, we detail how each type of segment is constructed. More details can be found in appendix.
@@
The instruction is a high-level description of what task is being performed in the current interaction sequence. The considered OpenAI Minecraft data includes meta information of each gameplay video, which depicts fundamental events that happened during in Minecraft gameplay, e.g. what block was just destroyed, what entity was just killed, what item was just crafted, etc. Such meta-information can provide a basic overview of what the player has been through in the gameplay. We therefore prompt an LLM into summarizing the gameplay with the meta information. The summary will be used as the instruction $D^{\text{inst}}_{t}$ of the current trajectory.
@@
The memory is the summary of what agents have finished in the previous interaction sequences. Due to the limited sequence length that the auto-regressive model can handle, the model needs to learn to summarize key information related to the task in historical interactions and ignore behaviors unrelated to instructions. The memory will be updated based on the results of each episode trunk and used for subsequent episode trunks. We therefore prompt an LLM into summarizing the gameplay with the meta information. The summary will then be used as the memory $D^{\text{mem}}_{t}$ of the current interaction trajectory. The memory prompt can be found in Appendix F.
@@
The thought is the agent’s reasoning and explanation of its own decisions. Previous methods have confirmed that using thought-enhanced interaction data helps language models understand decision-making [21]. Compared to labeling thoughts by humans [50, 6], we assume that thought is an intermediate variable that can be determined by the actions taken and observations made before and after the action, which is similar to an Inverse Dynamics Model [2]. We therefore prompt an LLM into estimating the thought of decisions with in-context learning, which will then be used as the thought $D^{\text{tht}}_{t}$ of the current behavior. Details can be found in Appendix E.
@@ 3.3 Architecture, Training, and Inference of OmniJARVIS
As illustrated in Figure 3, OmniJARVIS is built upon a pretrained MLM. We augment the original vocabulary of the MLM with additional tokens from the behavior tokenizer. Specifically, as we adopted the $[a,b,c]$ FSQ configuration (Section 2), we augment with $a+b+c$ new tokens as each behavior comprises $n$ behavior tokens $s^{bhv}_{1},\dots,s^{bhv}_{n}$ corresponding to $n$ FSQ levels. We formulate the learning objective of OmniJARVIS following [7, 37] in a prefix language modeling fashion. For a batch $\mathcal{B}$ of token sequence $s$, we optimize OmniJARVIS via:
@@ 4.1 Overview
The training of the OmniJARVIS is divided into two stages. In the first step, we use a self-supervised training method to train a Behavior Tokenizer, including the Encoder and Decoder jointly. We use FSQ as a quantization method and build a codebook with 8*8*8*6*5 discrete codes. The training data for Behavior Tokenizer comes from Contractor Dataset [2], which is a collection of Minecraft gameplay videos. The training parameters and details remain consistent with GROOT, which can be found in Appendix A.
In the second stage, we use this behavior tokenizer to process Minecraft offline trajectories to obtain behavior token sequences. We add 35 (8+8+8+6+5) additional tokens to the MLM tokenizer as behavior tokens for unified representation, so each time the VLA needs to output a continuous sequence of 5 tokens to represent a complete behavior. We use GPT-3.5 to synthesize thought, memory, and instruction to raw offline datasets to build complete interaction data. The specific prompt can be found in Appendix E. These data collectively constitute the embodied instruction-following dataset of OmniJARVIS, including 600k trajectories and about 900M tokens.
The training dataset of OmniJARVIS further includes a large amount of QA data about Minecraft. We generate a large number of seed questions about these texts using web pages on the Minecraft wiki. Then, we use the self-instruct method to generate a large number of creative questions and instructions. This constructed QA dataset consists of 300k conversations with about 90M tokens. During the training process, the QA data and instruction-following data are mixed, with a total of about 1T tokens, to train OmniJARVIS. In specific, we SFT (supervised finetune) LLaVA-7B [31]. The details can be found in Appendix A. To further demonstrate the generalizability of the method, we also fine-tune LLaVA at different scales and VLM Fuyu-8B with different architectures. The relevant results are presented in Section 4.5 and Section 4.6.
@@
We conduct experiments in the complex and open-world environment of Minecraft, a voxel-based 3D video game that has garnered significant attention from real-life research due to its popularity and diverse mechanics [18, 16]. We first evaluate OmniJARVIS with atomic tasks, which are skill-level tasks, testing VLAs’ ability to follow simple and straightforward instructions. Then we evaluate OmniJARVIS with programmatic tasks, which require the agent to obtain an item starting from an empty inventory. The success of these tasks requires VLAs to decompose the provided instruction into atomic-level subtasks, and hence tests VLAs’ complex reasoning ability. Finally, we test OmniJARVIS with open-world embodied question-answering benchmarks and creative free-form instruction-following. We also conduct ablation experiments of OmniJARVIS with different behavior tokenizers, different training dataset formats, and different vision tokenizations. Finally, we explore the generalization abilities of OmniJARVIS of Atari Games and the scaling potential of OmniJARVIS with different models and data scales.
4.2 Main Results I: Short-horizon Atomic Tasks